• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Lindgren

    Log entry time 16:51:59 on June27,2010

    Entry number 328757

    keyword=shift summary

    Day Shift Summary
    8:00 LHRS dipole P0 and Septum already set for delta scan = -2% when shift started. Shift started with no beam.
    9:30 Beam returned
    After beam returned we checked beam position with no target in.
    MCC changed beam position slightly to conform to position on whiteboard. After beam position was corrected we reset the attenuator and slits to get 10 na of beam.
    Need to keep event rate down to a about 2-3 K
    Took 500K events with optics target
    Changed delta scan to -4% and took 500K events with optics target
    CAUTION Septum magnet was not changed to match LHRS dipole
    Changed delta scan to -10%
    Increased beam current by opening slits to increase rate
    Took a few runs of 500K events with optics target. NMR lock a little unstable.
    CAUTION Septum magnet was not changed to match LHRS dipole
    The -4% and - 10% optics runs may not be useful for vertex reconstruction. Sorry guys.
    The label PREX P0 SET should be changed to APEX P0 SET to avoid similar mistakes. I know we should have known better.
    12:30 Call for Hall A Access to reverse polarity and repair NMR probe.
    15:30 Polarity reversal completed and dipole and quads recycled
    16:00 Still trying to repair NMR